# WARDEN: Multi-Directional Backdoor Watermarks for Embedding-as-a-Service Copyright Protection

### v1

@@ 4.1 Experimental Settings
To benchmark our attack and defense, we employ standard NLP datasets: Enron Metsis et al. (2006), SST2 Socher et al. (2013), AG News Zhang et al. (2015), and MIND Wu et al. (2020). We use the Enron dataset for email spam classification. AG News and MIND are news-based and used for recommendation and classification tasks. We use SST2 for sentiment classification. The statistics of these datasets are reported in Table 1.
@@
To evaluate our techniques, we adopt the following aspects of metrics:
  • (Downstream) Task Performance We construct a multi-layer perceptron (MLP) classifier with the EaaS embeddings as inputs. The quality of the embeddings is measured by the accuracy and $F_{1}$-score of the classifiers on the downstream task of predicting the class label.
  • (Reconstruction) Attack Performance We measure the closeness of reconstructed target embedding(s) (more details in Section 4.2) with original target embedding(s) by reporting cosine similarity.
@@ 4.2 CSE Experiments
CSE is designed to assist model extraction attack bypassing post-publish copyright verification. Hence, we evaluate whether we bypass the copyright verification using the same watermark detection metrics with an opposite objective, i.e., lower p-value and the absolute values of $\Delta$ metrics close to zero.
@@
Table 2: Copyright verification can be bypassed when the target direction is known and removed from the provided embeddings.
One of the critical elements in the EmbMarker is the secret target embedding ($t$) used for adding the watermark. The objective of CSE is to recover and erase this direction from the provided embeddings to circumvent copyright verification. We show later (see Table 3) that such elimination is feasible and does not deteriorate the EaaS quality. To demonstrate that, we start with a simplified case where we assume access to this target embedding and directly remove this direction. Expectedly, as seen in Table 2, we can bypass the copyright verification with minimal impact on the downstream utility performance. Moreover, this validates the importance of the projection technique employed in CSE. Additionally, this raises another technique’s vulnerability of ensuring target embedding is kept secure.
@@
In a successful attack, the principal components ${\bm{c}}^{\langle k\rangle}$ removed from the embeddings erase the watermark by recovering the target embedding. To validate this conjecture, we model and solve an optimization problem as defined in Equation 9 where a linear combination of ${\bm{c}}^{\langle k\rangle}$ results in the recovered target embedding ${\bm{w}}$. We then calculate cosine similarity to the target embedding ${\bm{w}}$. A high cosine similarity demonstrates the CSE technique’s effectiveness. For CSE, the reconstructed target embedding is extremely (99+cosine similarity) close to the original target embedding (more in following Section 4.2).
@@
In Table 3, CSE along with model extraction attack is proved effective in removing the influence from EmbMarker. Detection performance dropped to almost the original case, which indicates the EaaS provider will not be able to detect the imitation attack performed by the attacker. In addition, for SST2, MIND, AG News datasets, the downstream performance dropped 1-2%, which demonstrates the attack preserves the embeddings utility. We skip attack performance in the Table 3 as they all have full (almost) watermark reconstruction. However, we discuss this in Section 4.3 where we observe varying values due to the ineffectiveness of CSE attack against WARDEN defense.
@@ 4.3 WARDEN Experiments
Figure 5: Impact of the number of watermarks ($R$) in WARDEN for SST2 dataset.
@@
Figure 6: Impact of the number of watermarks ($R$) in WARDEN against CSE on SST2 dataset. Note: ‘Recon. Cos. Sim.’ (the Red line) represents the minimum reconstructed cosine similarity among all watermarks ${\bm{W}}$.
@@
Figure 7: Impact of GS extension on WARDEN for SST2 dataset. The observations are in line with normal WARDEN (Figure 6) results with the only difference being stronger metrics.
@@ 5 Conclusion
We test our WARDEN defense against CSE attack, acknowledging that various other attacks might overcome the uni-directional watermark approach. Nonetheless, we are of the view that the confusion created by employing multiple watermarks is likely to be effective against attacks similar to CSE attack. We also know that by having access to the original ground-truth watermarking vectors, combined with the GS process, one can eliminate the WARDEN watermarks, as discussed in Appendix D.3. However, it is the service provider’s responsibility to maintain the confidentiality of their watermarking keys.
Finally, regarding potential negative societal impacts, we develop a CSE attack, which could aid attackers in circumventing EaaS IP infringements. Yet, we argue that it is critical to show the possibility of such attacks and make users aware of them. The usual first step in security is to first expose the vulnerability. Additionally, to mitigate the aforementioned threat, we contribute an improved watermarking technique, WARDEN.
Acknowledgments
This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.
@@ C.2 Number of Clusters ($N$)
Figure 9: Impact of cluster numbers ($N$) in CSE for different datasets.
@@ C.3 Number of Principal Components ($K$)
Figure 11: Impact of number of principal components ($N$) in CSE for different datasets.
@@ C.4 Impact of Attacker Model Size
Table 6: Impact of extracted model size on CSE performance.
@@ D.1 Remaining Dataset Results
Figure 12: Impact of number of watermarks ($R$) in WARDEN for different datasets. As expected, detection performance (yellow and green lines) shows an upward trend with stable task performance.
@@
Figure 13: Impact of number of watermarks ($R$) in WARDEN against CSE for different datasets. The observation is similar to Figure 12, along with a decreasing trend in attack performance (red line) demonstrating the effectiveness of WARDEN defense against CSE attack.
@@
Figure 14: Impact of number of watermarks ($R$) in WARDEN GS extension for remaining datasets. Same trend as Figure 12, but stronger metrics.
@@
Figure 15: Impact of number of watermarks ($R$) in WARDEN GS extension against CSE for different datasets. Same trend as 13, but stronger metrics.
@@ D.2 Number of Principal Components ($K$) in CSE
Figure 16: Against WARDEN, impact of number of principal components ($K$) in CSE for different datasets.

### v2

@@ 4.1 Experimental Settings
To benchmark our attack and defense, we employ standard NLP datasets: Enron Metsis et al. (2006), SST2 Socher et al. (2013), AG News Zhang et al. (2015), and MIND Wu et al. (2020). We use Enron dataset for email spam classification. AG News and MIND are news-based and used for recommendation and classification tasks. We use SST2 for sentiment classification. The statistics of these datasets are reported in Table 1.
@@
To evaluate different aspects of our techniques, we adopt the following metrics:
  • (Downstream) Task Performance We construct a multi-layer perceptron (MLP) classifier with the EaaS embeddings as inputs. The quality of the embeddings is measured by the accuracy and $F_{1}$-score of the classifiers on the downstream tasks.
  • (Reconstruction) Attack Performance We measure the closeness of reconstructed target embedding(s) (more details in Section 4.2) with original target embedding(s) by reporting their cosine similarity.
@@ 4.2 CSE Experiments
CSE is designed to assist model extraction attack bypassing post-publish copyright verification. Hence, we evaluate whether we are able to bypass the copyright verification using the same watermark detection metrics with an opposite objective, i.e., lower p-value and the absolute values of $\Delta$ metrics close to zero.
@@
Table 2: Copyright verification can be bypassed when the target direction is known and eliminated from the provided embeddings.
One of the critical elements in the EmbMarker is the secret target embedding (${\bm{w}}$) used for adding the watermark. The objective of CSE is to recover and erase this direction from the provided embeddings to circumvent copyright verification. We show later (see Table 3) that such elimination is feasible and does not deteriorate the EaaS quality. To demonstrate that, we start with a simplified case where we assume access to this target embedding and directly remove this direction. Expectedly, as seen in Table 2, we can bypass the copyright verification with minimal impact on the downstream utility performance. Moreover, this validates the importance of the projection technique employed in CSE. Additionally, this raises another technique’s vulnerability of ensuring target embedding is kept secure.
@@
In a successful attack, the principal components ${\bm{c}}^{\langle k\rangle}$ removed from the embeddings erase the watermark by recovering the target embedding. To validate this conjecture, we model and solve an optimization problem as defined in Equation 9 where a linear combination of ${\bm{c}}^{\langle k\rangle}$ results in the recovered target embedding ${\bm{w}}$. We then calculate cosine similarity to the target embedding ${\bm{w}}$. A high cosine similarity demonstrates the CSE technique’s effectiveness. For CSE, the reconstructed target embedding is extremely (99+cosine similarity) close to the original target embedding (more in following Section 4.2),
@@
In Table 3, CSE along with model extraction attack is proved effective in removing the influence from EmbMarker. Detection performance dropped to almost the original case, which indicates the EaaS provider will not be able to detect the imitation attack performed by the attacker. In addition, for SST2, MIND, AG News datasets, the downstream performance dropped 1-2%, which demonstrates the attack preserves the embeddings utility. We skip attack performance in the Table 3 as they all have (almost) full watermark reconstruction. However, we discuss this in Section 4.3 where we observe varying values due to the ineffectiveness of CSE attack against WARDEN defense.
@@ 4.3 WARDEN Experiments
Figure 5: The impact of the number of watermarks ($R$) in WARDEN for SST2 dataset.
@@
Figure 6: The impact of the number of watermarks ($R$) in WARDEN against CSE on SST2 dataset. Note: ‘Recon. Cos. Sim.’ (the Red line) represents the minimum reconstructed cosine similarity among all possible watermarks in ${\bm{W}}$.
@@
Figure 7: The impact of GS extension on WARDEN for SST2 dataset. The observations are in line with normal WARDEN (Figure 6) results with the only difference being stronger metrics.
@@ 5 Conclusion
We test our WARDEN defense against CSE attack, acknowledging that various other attacks might overcome the uni-directional watermark approach. Although publishing the WARDEN algorithm to the public may inspire future attacks against it, we do not foresee it to be a trivial task, as the capability of WARDEN can be enhanced by using more conservative strategies, e.g., more stealthy trigger patterns and watermarking techniques.
We also know that by having access to the ground-truth watermarking vectors, combined with the GS process, one can eliminate the WARDEN watermarks, as discussed in Appendix D.3. However, it is the service provider’s responsibility to maintain the confidentiality of their watermarking keys.
We also note the false positive of the p-value for non-watermarked models when a large number of watermark vectors are augmented (Figure 18), even though other metrics rectify this incorrect signal. We suggest service providers conduct a preliminary study to select the number of watermarks for WARDEN. In this regard, the current work is an empirical observation study, and theoretical analysis might help decide the optimal number of watermarks. In future, we will endeavour to investigate advanced watermarking mechanisms focusing on defence purposes.
Social Impact
We developed a CSE attack, which could aid attackers in circumventing EaaS IP infringements. We agree that with CSE, any existing system using EmbMarker is vulnerable. Yet, we argue that it is critical to show the possibility of such attacks and make users aware of them. The usual first step in security is to first expose the vulnerability. Additionally, to mitigate the aforementioned threat, we contribute an improved watermarking technique, WARDEN, which could be incorporated with minimal effort.
Acknowledgements
We would like to appreciate the valuable feedback from all anonymous reviewers. This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative. Qiongkai Xu would like to express his gratitude to the FSE Staff Travel Scheme and FSE DDRI grant for the support in both travel and research. Anudeex Shetty would like to express his gratitude to Newman College for the support in travel and research.
@@ C.2 Number of Clusters ($n$)
Figure 9: The impact of cluster numbers ($n$) in CSE for different datasets.
@@ C.3 Number of Principal Components ($K$)
Figure 11: The impact of number of principal components ($N$) in CSE for different datasets.
@@ C.4 Impact of Attacker Model Size
Table 6: The impact of extracted model size on CSE performance.
@@ D.1 Remaining Dataset Results
Figure 12: The impact of number of watermarks ($R$) in WARDEN for different datasets. As expected, detection performance (yellow and green lines) shows an upward trend with stable task performance.
@@
Figure 13: The impact of number of watermarks ($R$) in WARDEN against CSE for different datasets. The observation is similar to Figure 12, along with a decreasing trend in attack performance (red line) demonstrating the effectiveness of WARDEN defense against CSE attack.
@@
Figure 14: The impact of number of watermarks ($R$) in WARDEN GS extension for remaining datasets. Same trend as Figure 12, but stronger metrics.
@@
Figure 15: The impact of number of watermarks ($R$) in WARDEN GS extension against CSE for different datasets. Same trend as 13, but stronger metrics.
@@ D.2 Number of Principal Components ($K$) in CSE
Figure 16: Against WARDEN, the impact of number of principal components ($K$) in CSE for different datasets.